# Comparing Hallucination Detection Metrics for Multilingual Generation

### v1

@@ 5 Human Evaluation of Multilingual Hallucination Metrics
R2-P. 0.21 30.05 89.08 - 0.19 53.28 46.25 -
RL-P. 0.17 28.54 85.35 - 0.13 50.31 49.93 -
NEO-P. 0.17 16.15 83.75 - 0.12 57.54 47.51 -
$\mathrm{DIFF}$ 0.34 56.11 94.14 75.00 0.31 65.85 60.90 63.19
$\mathrm{ENT}$ 0.49 65.32 94.96 78.97 0.35 68.00 63.6964.18
$\mathrm{CON}$ 0.08 31.56 87.49 69.77 -0.19 53.18 57.43 59.70
Table 5: Sentence- and atomic-fact-level verifiable hallucination detection in human evaluation. F denotes factual examples and NF denotes non-factual examples. The numbers in yellow have p-values of the correlation is larger than 0.05. Note that the correlation coefficient of $\mathrm{CON}$ should be expected as negative.

### v2

@@ Appendix D Generation Prompt Templates
random - 18.34 82.56 - 54.28 48.31
$\mathrm{DIFF}$ 0.28 49.11 92.41 0.29 63.93 62.49
$\mathrm{ENT}$ 0.31 45.43 90.27 0.33 66.67 63.69
$\mathrm{CON}$0.14 31.56 92.20 -0.21 60.18 56.90
$\mathrm{mFact}$ 0.31 31.70 85.52 0.28 64.84 60.10
$\mathrm{Seahorse}$ 0.08  29.28 85.80 0.09  60.94 57.44
Table 13: Sentence- and atomic-fact-level verifiable hallucination detection in human evaluation in English comparing automatic metrics and human support rate, with the best agreement with humans indicated in bold. F denotes factual examples and NF denotes non-factual examples. p-values of this correlation is larger than 0.05.